CHARACTERIZATION OF LINE BROADENING AND SHIFT PARAMETERS OF OZONE FOR SPEC-TROSCOPIC DATABASES

\underline{M. A. H. SMITH}, Atmospheric Sciences, NASA Langley Research Center, Mail Stop 401A, Hampton, VA 23681-2199.

Prior to 1997, most reported measurements of ozone line broadening and shift parameters involved the stronger transitions with 0 ≤ K′′_{a} ≤ 4. The broadening coefficients for this narrow range of K′′_{a} could be empirically modeled within the uncertainty of the measurements by polynomials in J′′^{a}^{b}. This allowed the prediction of broadening coefficients for transitions that had not been measured in the laboratory or calculated with detailed molecular interaction models. However, more extensive measurements involving higher K′′_{a} transitions in the ν_{1} and ν_{2} bands^{c}^{d} show that polynomials in J′′ do not adequately reproduce the measured broadening coefficients for the entire range of observed K′′_{a}. This paper reports a set of polynomials in J′′ and K′′_{a}, each valid for a specific range of K′′_{a} values, that reproduce most published measured air- and self-broadening coefficients of O_{3} within their experimental uncertainties. Dependence of air-induced line shift coefficients on the vibrational and rotational quantum numbers will also be examined.

aJ.-M. Flaud, C. Camy-Peyret, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and A. Goldman, Appl. Opt. \underline{29}, 3667-3671 (1990).

bM. A. H. Smith, C. P. Rinsland and V. Malathy Devi, J. Mol. Spectrosc. \underline{147}, 142-154 (1991).

cM. A. H. Smith, V. Malathy Devi, D. C. Benner and C. P. Rinsland, J. Mol. Spectrosc. \underline{182}, 239-259 (1997).

dV. Malathy Devi, D. C. Benner, M. A. H. Smith and C. P. Rinsland, J. Mol. Spectrosc. \underline{182}, 221-238 (1997).